Circular No. 8679

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 73P/SCHWASSMANN-WACHMANN Another companion to comet 73P (cf. IAUC 8659, 8663) has been

found by R. A. Tucker (Tucson, AZ, 0.35-m reflector, on Feb. 20.4 and 22.4 UT) and E. J. Christensen (at mag 17.2 with a 15′′ coma and a fanlike tail ≈ 20′′ long in p.a. 250^{o}–310^{o} on Mt. Lemmon Survey images taken on Feb. 24.5, with additional images also acquired on Feb. 25.5), denoted component ‘G’ on MPEC 2006-D47. Z. Sekanina, Jet Propulsion Laboratory, notes that this companion is not likely to be ‘E’ or ‘F’. The positions of component ‘G’ are consistent with the orbit for component ‘B’ (cf. MPEC 2006-B27, MPC 55982) but with T = 2006 June 8.11 TT (‘B’ has T = June 7.94, and ‘C’ has T = June 6.95; cf. MPC 55524).

COMET 41P/TUTTLE-GIACOBINI-KRESÁK This comet has been observed for the first time since 2001 via images

taken in the course of the LINEAR survey on Feb. 3.2 UT and in the course of the Mt. Lemmon Survey on Feb. 24.2 and 25.2 (cf. MPEC 2006-D45). E. J. Christensen, Lunar and Planetary Laboratory, reports that the Mt. Lemmon images on Feb. 24 show a coma of total mag 17.0 and size 15′′ × 20′′, elongated along the axis in p.a. 100^{o}/280^{o}.

COMETS C/2005 Y7, Y8, Y9, Y10 (SOHO) Additional comets have been found on SOHO website images (cf. IAUC

8672) — all being Kreutz sungrazers except for C/2005 Y8 (Meyer group), which was stellar in appearance but elongated horizontally (3×1 pixels), reaching mag 6.9 at 5.1R_{⊙} on Dec. 25.079 UT. C/2005 Y7 was small, stellar, and very faint in C3 images, and visible in C2 images as an extremely faint, fuzzy streak. C/2005 Y9 was stellar and condensed in C3 images, becoming teardrop-shaped near the C2 field-of-view, while in C2 images it showed a starlike head (reaching mag 4.0 at 7.9R_{⊙} on Dec. 27.412) with a thin, faint tail that reached a length of ∼ 28′ at 6.0R_{⊙} on Dec. 27.588. C/2005 Y10 was very faint, small, and slightly diffuse.

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 Y7 Dec. 24.779 18^{h}21.^{m}8 −26^{o}22′ C3/2 RK 2006-D03
C/2005 Y8 24.963 18 15.2 −22 39 C2 RK 2006-D03
C/2005 Y9  26.221 18 29.5 −28 45 C3/2 SF 2006-D03
C/2005 Y10  27.229 18 30.0 −24 57 C2  QY 2006-D03
2006 February 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green